there are 443 students in a graduating class. there are 103 more girls than boys. how many boys are there?
Answer by user_dude2008(1862)

b+g=443
g=b+103

b+g=443
b+b+103=443

2b=340

b=170 .... 170 boys

g=b+103=273 .... 273 girls